1) write the equation in slope intercept form (solved for y), when possible. Through (6,-7) parallel to the y-axis


2) write the equation in slope intercept form (solved for y), when possible. Through (-1,-6) perpendicular to the y-axis
3) Find an equation of the line with the given slope that passes through a given point. write the equation Ax+By=C .... m= -1, (-5,-9)

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

2) write the equation in slope intercept form (solved for y), when possible. Through (-1,-6) perpendicular to the y-axis
Plot the point and draw the line perpendicular to the y-axis.
It passes thru all the points where y = -6.
Equation: y = -6
==============================
3) Find an equation of the line with the given slope that passes through a given point. write the equation Ax+By=C .... m= -1, (-5,-9)
Use the form y = mx+b
-9 = -1*-5 + b
b = -14
----
Equation: y = -x-14
----
Change the form:
x + y = -14
===========================
Cheers,
Stan H.

Answer by solver91311(24713)



I already answered this one.

A line parallel to the -axis is a vertical line. Vertical lines have a couple of interesting characteristics. In the first place, ALL of the -coordinates of the set of ordered pairs that comprise the line have to be identical. Since all of the -coordinates are equal, no matter which two points you choose for the purposes of computing the slope, the slope fraction will have a zero denominator. Hence, the slope of any vertical line is undefined.

You can't write the equation of a vertical line in slope-intercept form because the slope quantity is undefined. However, since the -coordinates of all the points on the line are identical, the equation of a vertical line passing through the point is uniquely defined by the equation

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism